OpenStudy (anonymous):

Please help me? Finding first 3 terms?

OpenStudy (anonymous):

I tried all these numbers here :/

OpenStudy (anonymous):

2

OpenStudy (anonymous):

can you explain why? @shelbydancer

OpenStudy (anonymous):

Hard to see, is it \[c_n = (-1)^n+1\]

OpenStudy (anonymous):

yes @iambatman

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

You just have to plug in numbers for this.

OpenStudy (anonymous):

\[c_1 = (-1)^1 + 1 = ?\]

OpenStudy (anonymous):

Do the same with 2 and 3.

OpenStudy (anonymous):

@iambatman still stuck, none of the numbers work when I plug them in

OpenStudy (anonymous):

@perl can you help me with this?

OpenStudy (anonymous):

What do you mean, what is \[(-1)^1+1\] ?

OpenStudy (anonymous):

Anything to the power of 1 just means it's the number alone, so for example: \[2^1 = 2, ~~~4^1 = 4\] and if you have something such as \[2^2 = (2)(2)\] note that it's being multiplied.

OpenStudy (anonymous):

Oh ok. I got C, is that right? @iambatman

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

thank you